OpenStudy (z4k4r1y4):

(a) specify a suitable solution that will reduce a differential equation of the form dy/dx = f(x/y) to a separable equation. derive an expression for dy/dx using your substitution.

OpenStudy (z4k4r1y4):

(b) apply this technique to find the general solution for: \[\frac{ dy }{ dx}=\frac{ 3x^2+8y^2 }{ 4xy}\] giving your answer in the form y^2=g(x)

OpenStudy (christos):

You could solve the expression with respect to y. And because y is to the power of 2 , you will get two solutions , a positive and a negative one. Then you could integrate those. Or you could approach this as an inverse implicit differentiation.

OpenStudy (z4k4r1y4):

@ganeshie8 any suggestions for part (a)?

ganeshie8 (ganeshie8):

Let \(y=vx\)

ganeshie8 (ganeshie8):

plugging that in the right hand side, \(f(x/y)\), eliminates \(y\) completely : \[f(\frac{x}{y}) = f(\frac{x}{vx}) = f(1/v)\]

ganeshie8 (ganeshie8):

\(y=vx\) differentiating both sides with respect to \(x\), we get : \(\dfrac{dy}{dx} = \dfrac{dv}{dx}*x+v*1\)

ganeshie8 (ganeshie8):

In the given differential equation, you can replace left hand side derivative with above expression

ganeshie8 (ganeshie8):

\(\dfrac{dy}{dx} = f(\frac{x}{y})\) changes to \( \dfrac{dv}{dx}*x+v*1 = f(1/v)\) which is separabel

OpenStudy (z4k4r1y4):

yes, i understand what you've done. thank you, that was very helpful.